Part A Problem 24.3 What is the potential difference between the plates? A parallel-plate air capacitor with a capacitance of 244 pF has a charge of magnitude 0.137 μC on each plate. The plates have a separation of 0.265 mm 0.5615 10 Submit My Answers Give Up Incorrect; Try Again; 3 attempts remaining Part B What is the area of each plate? Use 8.85x10-12 F/m for the permittivity of free space. Submit My Answers Give Up Part C What is the electric field magnitude between the plates? VIm Submit My Answers Give Up Part D What is the surface-charge density on each plate?
Answer #1

Here:

(a) potential diff = Q / C =   0.137 x 10-6 / 244 x10-12 =    561.48 volts

(b)   C = εA /d     

so        

A = d C / ε = 0.000265 * 244 x 10-12 / 8.85 x10-12 =   0.007303m2

(c) E = V / d = 561.48 /0.000265 = 2118792.5 V/m

(d) σ = Q / A = 0.137 x10-6 / 0.007303 =   1.876 x 10-5 C/m2